Common Stock value: Constant Growth Seagate technology is a global leader in data storage solutions and a high-yield dividend player. From 2015-2019, Seagate paid to following per-share dividends:

2019-$2.57

2018-$2.26

2017-$1.86

2016-$1.19

2015-$1.57

Assume that the historical annual growth rate of Seagate dividends is an accurate estimate of the future constant annual dividend growth rate. Use 22% required rate of return to find the value of Seagate's stock immediately after it paid its 2019 dividend of $2.57.

The value of Seagate's stock is?
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

co Year 2019 2018 2017 2016 2015 - moroFM Dividend $2.57 $2.26 $1.86 $1.19 $1.57 Growth rate 13.72% 21.51% 56.30%. -24.20% Av

Cell reference -

_ A B D Year Dividend 2019 2018 2017 2016 2015 2.57 2.26 1.86 1.19 Growth rate =C3/C4-1 =C4/C5-1 =C5/C6-1 =C6/07-1 1.57 Avera

Hope it will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.